OpenStudy (anonymous):

what number must you add to complete the square? x^2+14x=-5

OpenStudy (anonymous):

iis 44 huh

OpenStudy (anonymous):

NOPE\[x^2+14x=-5 \]\[x^2 + 14x + (\frac{ 14 }{ 2 })^2 = -5 + (\frac{ 14 }{ 2 })^2\]\[x^2 + 14x + 49 = -5 + 49\] you would add 49 to both sides

OpenStudy (anonymous):

Whenever you want to solve this type of problem: